OpenStudy (anonymous):

For all r≠±2, r2−5r+6 (line in the middle r2−4 = ?

OpenStudy (anonymous):

rational expressions

OpenStudy (anonymous):

are we just simplifying?

OpenStudy (anonymous):

yes just simplifying

OpenStudy (anonymous):

okay then we first factor the numberator

OpenStudy (anonymous):

numerator*

OpenStudy (anonymous):

r^2-5r+6=(r-2)(r-3)

OpenStudy (anonymous):

The bottom factors as a difference of square: r^2-4=(r-2)(r+2)

OpenStudy (anonymous):

oh no i'm sorry we're solving for r≠±2

OpenStudy (anonymous):

okay thats fine, still follow the steps i have showed you

OpenStudy (anonymous):

So, now we have : (r-2)(r-3) -------- (r-2)(r+2)

OpenStudy (anonymous):

now, we diivde out (r-2)

OpenStudy (anonymous):

and we are left with out solution: (r-3) ---- (r+2)

OpenStudy (anonymous):

are we good?

OpenStudy (anonymous):

no I haven't taken math in so long, i'm so lost, can you possibly type it in one text step by step if possible

OpenStudy (anonymous):

ok i think i got it, just had to go over it few times :) thanks a lot Lagrangeson678